### v2

# Dynamical dark energy in the light of DESI 2024 data

Abstract
The latest findings from the DESI (Dark Energy Spectroscopic Instrument) data release 1 (DR1) [3], combined with data from the cosmic microwave background and supernovae, suggest a preference for dynamical dark energy over the cosmological constant. This study has considered the Chevallier-Polarski-Linder (CPL) parameterization for the dark energy equation of state (EoS) and has indicated a possible phantom barrier crossing in the recent past. Despite CPL being the most commonly used parameterization, recent research has pointed out issues with its prior selection and parameter degeneracies. In this paper, we introduce an alternative two-parameter parameterization of the dark-energy EoS, at high redshifts, this new parameterization can be approximated to the CPL form. Our findings also indicate that the current value of the EoS of dark energy resembles quintessence, with evidence of a recent crossing of the phantom barrier, supporting the conclusions in  [3]. Furthermore, our model significantly reduces the Hubble tension to about $2.8\sigma$ when compared to Hubble Space Telescope and SH0ES data  [51], and to $1.6\sigma$ with standardized TRGB and Type Ia supernova data  [56]. Bayesian model selection using Bayes factors shows a moderate preference for our parameterization over the $\Lambda$ CDM model, aligned with the DESI2024 results and favoring dynamical dark energy.
journal: Physics Letters B
[…]
[]NAS, Centre for Theoretical Physics & Natural Philosophy, Mahidol University, Nakhonsawan Campus, Phayuha Khiri, Nakhonsawan 60130, Thailand
1 Introduction
A wide variety of cosmological observations  [50, 44, 13, 6] have indicated that our universe is currently undergoing a mysterious accelerated expansion phase. No known form of matter can account for this accelerated expansion. Although the present observations align with the cosmological constant  [41] as the driving force behind the accelerated expansion, significant theoretical issues have been encountered over time, such as the cosmological constant problem, the coincidence problem, etc., and an increasing number of observational challenges.
Recently, as our precision in the measurement of the cosmological parameters has improved significantly, the estimation of different cosmological parameters from different cosmological observations shows a discrepancy between their estimation. The most significant one is the Hubble tension. A statistically significant discrepancy of the order of $\simeq 5.3\sigma$ [49] has been reported from the early universe measurements and the late time distance ladder measurement of the current value of the Hubble parameter ($H_{0}$). The early universe measurements like CMB from Planck collaboration  [4], BAO  [7, 16, 1, 40], and BBN  [8] data estimated the current value of the Hubble constant to be $H_{0}\sim(67.0-68.5)$ km/s/Mpc by considering the cosmological constant ($\Lambda$) as the candidate for dark energy. Whereas the measurement of $H_{0}$ using distance ladder measurements such as SH0ES  [48] and the collaborations of H0LiCOW  [65] reported that its value was $H_{0}=(74.03\pm 1.42)$ km / s / Mpc. This discrepancy indicates that the dynamics of the dark energy might be richer than simply a constant.
Considering dark energy as dynamical is one approach to alleviating the challenges posed by the cosmological constant. A variety of dynamical dark energy models have been suggested, each presenting unique benefits and limitations [11, 14]. Despite more than two decades since the discovery of the accelerated expansion of the universe and extensive efforts to explain it, the underlying physics remains unresolved, with no clear consensus. Attempts have been made to reconstruct various cosmological parameters by parameterizing them and validating them against observational data. The parameterization technique is crucial as it allows us to model the unknown physics of dark energy effectively without altering the underlying theory of gravity. In this approach, a functional form is assumed for certain cosmological parameters, such as the equation of state (EoS) of dark energy, the Hubble parameter, or the dark energy density, and the model is constrained using state-of-the-art cosmological data.
The simplest dynamical dark energy model is the fluid description of the dark energy in which the dark energy equation of state is parameterized phenomenologically and the most popular parameterization of the dark energy EOS is the CPL parameterization or $w_{0}w_{a}$ CDM model in which the EOS is parameterized as $w(a)=w_{0}+w_{a}(1-\frac{a}{a_{0}})$. Recent DESI 2024 Data Release 1 (DR1)  [3] has constrained the CPL model and found the preference of more than $2\sigma$ for the dynamical dark energy over $\Lambda CDM$. Studies have reported similar evidence for the CPL model  [63, 42]. On the other hand  [24] reported that the best model to fit the data is not CPL when comparing a wide number of dynamical dark energy models. In  [27, 59], the authors have shown concern about the choice of prior and degeneracy, which can affect the result reported in  [3]. DESI 2024 data have also been used to constrain other cosmological models [39, 18, 22, 66, 64, 33, 15, 26] showing a preference for dynamical dark energy.
In this study, we propose an alternative parameterization of the dark energy equation of state (EoS) inspired by the dynamics of the quintom scalar field model [53], which includes contributions from both quintessence and phantom scalar field. An interesting feature of this proposed parameterization is that, at a high redshift, the EoS takes the form of CPL parameterization, while differing at a low redshift. This suggests that our model may exhibit richer behavior than the CPL parameterization in the low redshift regime. We utilized the Pantheon Plus data with the SH0ES prior, along with the DESI BAO 2024 data and a compressed Planck likelihood, to constrain our model. Our results were compared with the $\Lambda$CDM model, and we applied Bayesian model selection using the Bayes factor to assess the preference for our model. In addition, we reported on the status of the Hubble tension in this context. To evaluate the performance of the proposed parameterization at the linear perturbation level, we computed the temperature anisotropies and the matter power spectrum (MPS), comparing them with those of the $\Lambda$CDM model.
The current work is presented in the following way: In Section 2, we provide an overview of the mathematical formulation and discuss about the proposed model. Section 3 discusses the observational data used in this work together with the constraints obtained on the cosmological parameters and the model comparison. Finally, we conclude our results and findings in Section 4.
2 Mathematical Setup and the Model
In a spatially flat homogeneous and isotropic universe filled with barotropic fluids, the Friedmann constrain equation can be written as
$\begin{split}\frac{H^{2}(z)}{H_{0}^{2}}&=\Omega_{\mathrm{m},0}(1+z)^{3}+\Omega_{\mathrm{r},0}(1+z)^{4}+\left(1-\Omega_{\mathrm{m},0}-\Omega_{\mathrm{r},0}\right)f_{\mathrm{DE}}(z),\end{split}$ (1)
where $\Omega_{m,0}$ and $\Omega_{r,0}$ are the current matter and radiation density respectively and $f_{\mathrm{DE}}(z)\equiv\rho_{\mathrm{DE}}(z)/\rho_{\mathrm{DE},0}$ is the normalized dark energy density given by the following expression,
$f_{\mathrm{DE}}(z)\equiv\frac{\rho_{\mathrm{DE}}(z)}{\rho_{\mathrm{DE},0}}=\exp\left(3\int_{0}^{z}[1+w(\tilde{z})]\mathrm{d}\ln(1+\tilde{z})\right).$ (2)
Although considering any candidate for dark energy the expansion of the universe is affected by the evolution of this normalized dark energy density $f_{\mathrm{DE}}(z)$. Due to the unknown nature of dark energy, the most popular and simple way to investigate the dynamics of dark energy is to consider a parameterization of the dark energy equation of state. In this approach, there is no need to alter the underlying theory of gravity, and all the important cosmological parameters can be successfully reconstructed. Here we propose a new two-parameter parameterization of the dark energy equation of state in the following form,
$\begin{split}w(a)=&{w_{a}}\left(\frac{a}{a_{0}}\right)Cos\left(\frac{a}{a_{0}}\right)\\ &-{w_{b}}\left(\frac{a}{a_{0}}\right)Cosh\left(\frac{a}{a_{0}}\right)-1\end{split}$ (3)
The reason behind introducing the trigonometric and hyperbolic function in the equation of state is inspired by the equation of state of the scalar field dark energy models of quintessence and phantom fields. In  [54, 62, 52] it has been shown the equation of state of a quintessence field ($w_{DE}>-1$) can be written as a cosine function whereas in  [38] it has been shown the equation of state of a phantom scalar field ($w_{DE}<-1$) can be written as the function of the cosh and for the quintom model it to be a function of the both cosine and cosh  [53]. The quintom models are known for its ability to cross the phantom divider line (PDL). Recent studies using model-independent approaches and data-driven approaches  [32, 35, 43, 55, 10, 29, 31, 58] have shown the indication of PDL crossing for the dark energy sector. Also results from DESI (2024)  [3] have also shown a preference for dynamical dark energy with a phantom barrier crossing. These results hint towards more complex late-time dynamics of the dark energy.
From the motivation above, we propose the current parameterization of the EoS of the dark energy. It consists of two dedicated parts: one corresponds to quintessence and the other to the phantom nature of the EoS. In cases $w_{b}=0$ and $w_{a}>1$ this EoS can show only quintessence-like behaviour since $0\leq Cos\left(\frac{a}{a_{0}}\right)\leq 1$ since $0\leq\frac{a}{a_{0}}\leq 1$. On the other hand, it can show only phantom-like behaviour when $w_{a}=0$ and $w_{b}>1$ since $Cosh\left(\frac{a}{a_{0}}\right)\geq 1$. The existence of both these terms in EoS expected to give rise to quintom-like behaviour with the possibility of smooth phantom barrier crossing. Another interesting nature of this proposed EoS is that at $\frac{a}{a_{0}}<<1$ this parameterization reduces to the CPL parameterization as follows.
$w(a)=\tilde{w_{0}}+\tilde{w_{a}}(1-\frac{a}{a_{0}}),$ (4)
where $\tilde{w_{0}}=(w_{a}-w_{b}-1)$ and $\tilde{w_{a}}={w_{b}}-{w_{a}}$. Due to this reason, we can expect more reacher behaviour of the dark energy at the low redshift than the CPL model.
We have implemented this proposed parameterization to the publicly available CLASS Boltzmann code [36, 37] to numerically understand its behavior and to constrain the model parameters against different cosmological observations we have used the MontePython MCMC cosmological code  [19].
3 Observational Data
To constraint the cosmological parameters, we have used the following data sets;
3.1 Pantheon Plus SH0ES
Type Ia supernovae are widely used as standard candles because of their relatively uniform absolute luminosity [46, 50]. In this analysis, we used the Pantheon Plus compilation sample of SN-Ia data  [20, 51] together with a SH0ES prior.
3.2 DESI BAO
The density of visible baryonic matter exhibits recurring, periodic fluctuations called baryon acoustic oscillations. These oscillations are essential standard rulers for precise distance measurements in cosmology. In this study, we have used the 2024 BAO observation data from the Dark Energy Spectroscopic Instrument (DESI) as noted in reference  [3]. Covering a redshift range of $z\in[0.1,4.2]$, the survey is divided into seven redshift bins. BAO provides measurements of the effective distance along the line of sight as
$\frac{D_{H}(z)}{r_{d}}=\frac{cr_{d}^{-1}}{H(z)}$ (5)
and along the transverse line of sight as,
$\frac{D_{M}(z)}{r_{d}}\equiv\frac{c}{r_{d}}\int_{0}^{z}\frac{d\tilde{z}}{H(\tilde{z})}=\frac{c}{H_{0}r_{d}}\int_{0}^{z}\frac{d\tilde{z}}{h(\tilde{z})}.$ (6)
The angle average distance is measured as,
$\frac{D_{V}(z)}{r_{d}}=\left[\frac{czr_{d}^{-3}d_{L}^{2}(z)}{H(z)(1+z)^{2}}\right]^{\frac{1}{3}}.$ (7)
[…]
3.3 Compressed Planck likelihood
In this analysis, we have used the Planck compressed likelihood following the approach in  [12]. While having limited computational resources this compressed likelihood is useful to obtain the same constraining accuracy as full Planck. The compressed likelihood uses the baryon physical density $\omega_{b}=\Omega_{b}h^{2}$ and the two shift parameters $\theta_{*}=r_{s}\left(z_{\text{dec }}\right)/D_{A}\left(z_{\text{dec }}\right),\quad\mathcal{R}=\sqrt{\Omega_{M}H_{0}^{2}}D_{A}\left(z_{\text{dec }}\right),$ where, the redshift decoupling is $z_{dec}$, and $D_{A}$ is the comoving angular diameter distance. Here we chose the mean values of the above parameters to be $100\omega_{b}=2.237\pm 0.015,100\theta_{s}=1.0411\pm 0.00031$, and $\mathcal{R}=1.74998\pm 0.004$, and the correlation matrix given in Appendix A of  [12].
3.4 Observational Constraints
We have considered flat priors on the cosmological parameters, $100~{}\omega_{b}:[0,4.6],\omega_{cdm}:[0.095,0.145]$. For the model parameters, we have considered the prior to be $w_{a}:[-3,1]$ and $w_{b}:[-1,1]$. The mean value and the corresponding $68\%$CL constraint obtained on the cosmological parameters of our analysis are given in Table1. The result is shown for the combined data sets of SN-Ia Pantheon Plus compilation [57] with SH0ES prior, BAO DESI2024  [3], together with the compressed Planck likelihood [12]. For comparison, the constraint on the $\Lambda CDM$ model and the proposed model as $wCDM$ are given together.
The 2D and 1D triangular plots of the cosmological parameters $100w_{b},w_{cdm},H_{0},\Omega_{DE}$ are shown in FIG.1. The posteriors for $\Lambda CDM$ are shown in blue and the ones for our $wCDM$ model are shown in red. It can be seen that the proposed model suggests a higher value of $H_{0}$ and $\Omega_{DE}$ compared to $\Lambda CDM$. In FIG.2 we have shown the 2D posterior distribution of $H_{0}$ versus $\Omega_{m}$ together with $\Lambda CDM$ for comparison. The deep and light green bands show, respectively, the constraints $1\sigma$ and $3\sigma$ on $H_{0}$ from  [47]. To quantify the status of Hubble tension in the proposed model we have used the estimator proposed in [23], $T_{H0}=\frac{\left|H_{0}-H_{0}^{loc}\right|}{\sqrt{\sigma_{H0}^{2}+\sigma_{\mathrm{loc}}^{2}}}$, where $T_{H0}$ denotes the estimation of the amount of tension, $H_{0}$ and $H_{0}^{loc}$ is the mean value of the Hubble parameter obtained from the current analysis and local universe observations, respectively, while $\sigma^{2}$ represents the corresponding variance of the posteriors. By comparing our results with the result obtained in  [51] $\left(H_{0}=73.04\pm 1.04\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}\right)$, from the Hubble Space Telescope and the SH0ES data, the tension status is $T_{H0}\simeq$ $2.86\sigma$. On the other hand, when comparing it with the reported value of $\left(H_{0}=73.22\pm\right.$ $2.06\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$ ) from the standardized TRGB and Type Ia supernova data sets in  [56], it reduces to $T_{H0}\simeq 1.69\sigma$. If we consider the classification of the tension in Table IV in  [23] for the first case, the tension is moderate and weak for the latter case.
Figure 3 shows the contour plot of posterior distributions of the model parameters $w_{a},w_{b}$ and the current value of the dark energy EOS $w_{DE}$. The current value of the EOS is $w_{DE}=-0.784_{-0.089}^{+0.10}$ which is quintessence in nature. To understand the evolution of the EOS of the dark energy in FIG.4 we have plotted the posterior probability $Pr(w|z)$ of the $w_{DE}$ against the redshift $z$. The deeper and lighter blue region shows the contour level $1\sigma$ and $2\sigma$, while the solid blue line shows the mean value. However, it can be seen that the current value of the EOS is in the quintessence region, but in the recent past, it has a phantom barrier crossing around $z\approx 0.35$ from phantom to quintessence. This finding of ours matches the one reported in  [3, 21]. Although we have obtained constraints on the main cosmological parameters as expected, it is necessary to verify the evolution of the Hubble parameter against the observed data. Therefore, we have plotted the posterior probability of $H(z)/(1+z)$ vs. $z$ in FIG. 5, along with observational data points from the SH0ES survey  [47] and Baryon Acoustic Oscillation (BAO) surveys  [9, 67, 17, 28]. The evolution of $H(z)/(1+z)$ shows the expected behavior of the proposed model.
[…]
At the bottom pan of FIG.6 and FIG. 7 we have shown the relative differences in $\Delta D_{l}=\left(D_{l}-D_{l}^{\Lambda CDM}\right)/D_{l}^{\Lambda CDM}$ and $\Delta P(k)=\left(P(k)-P(k)^{\Lambda CDM}\right)/P(k)^{\Lambda CDM}$ to compare with the $\Lambda CDM$ model. For $D_{l}$, the proposed model shows a small deviation from $\Lambda CDM$ at higher multipoles, and for $P(k)$ it can be seen at the higher scale. The PPF approximation (Parameterized Post-Friedmann) [30] has been used to calculate $D_{l}$ and $P(k)$. PPF approximation is useful for perturbations to cross the phantom divider smoothly. This approximation has already been implemented in the CLASS code by default. Both the plots in FIG.6 and in FIG.7 suggest the consistency of the proposed model not only at the background level but also at the perturbation level.
3.5 Comparison with $\Lambda CDM$
We use the Bayes factor concept to determine the preference of this model over the $\Lambda CDM$ model. Bayes factor is calculated as $\ln B_{w\Lambda}=\ln\mathcal{Z}{w}-\ln\mathcal{Z}\Lambda$, where $\mathcal{Z}$ represents Bayesian evidence and the suffixes $w$ and $\Lambda$ represent the current $wCDM$ and $\Lambda$CDM models, respectively.
In general, Jeffrey’s scale is used to find the preference of one model over another model. According to this scale, a negative preference will be considered if $\lvert\ln B_{w\Lambda}\rvert<1$, on the other hand positive, moderate, and strong preferences will be considered if $\lvert\ln B_{I\Lambda}\rvert>1$, $\lvert\ln B_{I\Lambda}\rvert>2.5$, and $\lvert\ln B_{I\Lambda}\rvert>5.0$, respectively  [60]. We have used a publicly available code MCEvidence [34] to directly calculate the Bayes factor from the MCMC chains. From Table.1 it can be seen that the evidence for our model is $4.77$ indicating moderate evidence for the current model over the $\Lambda$CDM model. The preference we find here using the Bayes factor concept is very similar to the evidence obtained for the CPL model in [3]. We also calculated $\delta\xi^{2}_{min}$ for our model in comparison with $\Lambda CDM$, obtaining the value of $-6.11$, which indicates a better fit to data for our proposed model.
Parameter $\Lambda$ CDM $w$CDM
$10^{-2}\omega_{b}$ $2.267^{+0.013}_{-0.014}$ $2.244^{+0.017}_{-0.012}$
$\omega_{cdm}$ $0.116^{+0.00095}_{-0.00071}$ $0.119^{+0.00077}_{-0.0013}$
$H_{0}$ $68.73^{+0.32}_{-0.47}$ $69.91_{-1.2}^{+1.1}$
$\Omega_{DE}$ $0.6917_{-0.0062}^{+0.0065}$ $0.7056^{+0.0061}_{-0.0052}$
$\Omega_{m}$ $0.2967^{+0.0055}_{-0.0044}$ $0.2943^{+0.0052}_{-0.0061}$
$w_{a}$ $-1.27^{+0.45}_{-0.37}$
$w_{b}$ $-0.59\pm 0.20$
$w_{de}$1 $-0.784_{-0.089}^{+0.10}$
$\delta{\chi_{min}^{2}}$ 0 $-6.11$
$\lvert\ln B_{w\Lambda}\rvert$ $4.77$
Table 1: The constraint ($68\%$CL) obtained on cosmological parameters from the combining data sets of Pantheon supernova survey with SH0ES prior  [57], DESI2024 data [3] together with the compressed Planck likelihood.
4 Conclusions
Recent findings from DESI2024 [3] show a preference for the dynamical nature of dark energy, including a potential transition from phantom to quintessence phases in the recent past. In this study, we introduce an alternative parameterization of the dark energy equation of state, motivated by quintom-type scalar field models, yielding results comparable to those reported in DESI2024 [3]. One interesting feature of our proposed parameterization is that, while it differs from the CPL model at low redshift, it reduces to the CPL form at high redshift. This means that it could provide a richer understanding of cosmology at lower redshifts compared to that of CPL. The model uses trigonometric cosine and hyperbolic cosine functions to represent the equation of state (EoS) for quintessence and phantom nature, respectively.
We used recently released BAO data from [3], Pantheon+ supernova data with SH0ES prior, and a compressed Planck likelihood to constrain the cosmological parameters. Our findings also suggest that the current equation of state (EoS) of dark energy is of the quintessence type, with a possible phantom barrier crossing in the recent past, which aligns with the results of DESI2024[3, 21]. The transition from phantom to quintessence occurs at approximately $z\simeq 0.35$. For this model, we also find that the current value of the Hubble parameter is higher than that in both the CPL and $\Lambda$CDM models. The Hubble tension is significantly reduced in our model, weakening to approximately $2.8\sigma$ compared to the Hubble Space Telescope and SH0ES data [51], and to $1.6\sigma$ when compared with standardized TRGB and Type Ia supernova data [56]. To evaluate its performance at the linear perturbation level, we plotted the temperature anisotropies and the matter power spectrum (MPS) and compared them to the $\Lambda$CDM model. A slight deviation from the $\Lambda$CDM model was observed at high multipoles and large scales, indicating that our model remains consistent even at the perturbation level. We also calculated the Bayes factor to assess the preference of this model over the $\Lambda$CDM model and found a moderate preference for our model, indicating the dynamical nature of dark energy over the cosmological constant.
5 Acknowledgement
The author acknowledges the use of the Chalawan High Performance Computing cluster, operated and maintained by the National Astronomical Research Institute of Thailand (NARIT).
Figure 1: A triangular plot illustrates the constraints on cosmological parameters for the proposed model $wCDM$ (in red), with the $\Lambda$CDM model (in blue) and CPL model (in green) for comparison.
Figure 2: The figure shows two-dimensional posterior plots on $(\Omega_{0},H_{0})$ plane. The value of the $H_{0}$ from SH0ES collaboration [47] is shown by the green band. The deep and light green regions represent $1\sigma$ and $3\sigma$ constraints respectively.
Figure 3: A triangular posterior plot showing the constraint obtained on the model parameters $w_{a},w_{b}$ and the dark energy EOS ($w_{DE}$).
Figure 4: This figure shows the posterior probability $Pr(w|z)$ of the dark energy equation of state (EOS). The dark blue area indicates the $1\sigma$ confidence level, while the light blue areas show the $2\sigma$ confidence level.
Figure 5: This figure shows $Pr(H(z)/(1+z))$ versus $z$ for the current model, compared with Sh0ES survey data  [47] and BAO surveys  [9, 67, 17, 28]. The dark blue area represents the $1\sigma$ confidence level, and the light blue areas represent the $2\sigma$ confidence level.
Figure 6: Plot of the CMB anisotropies for the $wCDM$ model (in yellow) using the mean value of the model parameters $w_{a},w_{b}$ together with the $\Lambda CDM$ model (in blue). The bottom panel shows the relative difference between the models $\triangle D_{l}=\left(D_{l}-D_{i}^{\Lambda CDM}\right)/D_{l}^{\Lambda CDM}$.
Figure 7: The Plot of the matter power spectrum for the $wCDM$ model (in yellow) using the mean value of the model parameters $w_{a},w_{b}$ together with the $\Lambda CDM$ model (in blue). The bottom panel shows the relative difference between the models $\triangle P(k)=\left(P(k)-P(k)^{\Lambda CDM}\right)/P(k)^{\Lambda CDM}$. Observations data points have been also shown for comparison.